OpenStudy (mikezack123):

If f(x) = x + 3, find the value of x if f(x) = 9.

OpenStudy (anonymous):

if f(x)=x+3 and f(x)=9, then, x+3=9

OpenStudy (anonymous):

The rest should be straightforward from there :)

OpenStudy (mikezack123):

so x= 6

OpenStudy (mikezack123):

Thanks :)